I don't understand how to put words into equations. For example, Martha sold 4 times as many tickets to the Chili Supper as Harvey sold. Together they sold 185 tickets. How many tickets did each person sell?

Answer by Maths68(1474)

let
Harvey sold = x tickets
Martha sold 4 times of x= 4x tickets
Together they sold = 185 tickets
Harvey sold + Martha sold = 185
x+4x=185
5x=185
5x/5=185/5
Divide by 5 both sides of above equation
x=37
So
Harvey sold = 37 tickets
Martha sold 4 times of x= 4*37 tickets
Martha sold 4 times of x= 148 tickets